If a full wave rectifier circuit is operating from 50Hz mains, the fundamental frequency in the ripple will be


        (a)    100 Hz


        (b)    25 Hz


        (c)    50 Hz


        (d)    70.7 Hz


Grade:10

3 Answers

Kevin Nash
10 years ago

(a)

In case of full wave rectifier,

Fundamental frequency = 2 × mains frequency

=      2 × 50 = 100Hz.

Chetan Patil
10 years ago

For full wave rectifier ripple frequency is given by 2Fn. Ripple fre. is 100 Hz

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

In case of full wave rectifier,
Fundamental frequency = 2 × mains frequency
= 2 × 50 = 100Hz.

Thanks and Regards